2.76g of Ag on being highly heated yields a residue weighing – 2Ag+CO2+1/2O2

Grade:9

1 Answers

Arun
6 years ago
Unlike other metal carbonates that usually decomposes into metal oxides liberating carbon dioxide, silver carbonate on heating decomposes into elemental silver liberating mixture of carbon dioxide and oxygen gas as Ag2CO3 (s) 2Ag (s) + CO2 (g) + 1/2 O2 (g) MW = 276g 2 × 108 = 216 g Hence, 2.76 g of Ag2CO3 on heating will give 216/276× 2.76 = 2.16 g Ag as residue